OpenStudy (patrickcharran97):

HELP I SUCK USING SUBSTITUTION! solve using substitution: -4x-6y=-6 -8x-6y=-18

jimthompson5910 (jim_thompson5910):

Start with equation 1 and solve for -6y -4x-6y=-6 -6y=-6+4x ------------------------------------------------------- -8x-6y=-18 -8x+(-6y)=-18 -8x+(-6+4x)=-18 ... replace -6y with -6+4x -8x-6+4x=-18 -4x - 6 = -18 I'll let you finish

jimthompson5910 (jim_thompson5910):

We are able to replace -6y with -6 + 4x since -6y = -6 + 4x

OpenStudy (patrickcharran97):

is it (3,-1)?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Notice both equations have \(-6y\) as terms, so it should be particularly easy to substitute for.$$-4x-6y=-6\\-6y=4x-6$$... so we may substitute into our other equation \(-8x-6y=-18\) to yield $$-8x+4x-6=-18\\-4x-6=-18\\-4x=-12\\x=3$$which we may plug into our earlier expression to determine \(y\):$$-6y=4x-6=4(3)-6\\-6y=6\\y=-1$$